The brick wall (of thermal conductivity 0.86 W/m ·◦ C) of a building has dimensions of 3.7m by 16m and is13cm thick.

How much heat flows through the wall in a 17.4 h period when the average inside and outside temperatures are, respectively, 9 ◦C and 8◦C?

Answer in units of MJ.


1
Expert's answer
2021-04-29T10:42:09-0400
"E=Pt=\frac{kA\Delta T}{L}t""E=\frac{(0.86)(3.7)(16)(9-8)}{0.13}(17.4\cdot 3600)=14.5\cdot 10^6 J""E=14.5\ MJ"